OpenStudy (anonymous):

Principle Amt: $5000, 8% compounded continuously; How many years would it take for the investment to grow to $7000?

OpenStudy (anonymous):

Anybody know the formula?

OpenStudy (anonymous):

A = P e^ (rt)

OpenStudy (anonymous):

ok, now we have A=5000^(.08*t) ?

OpenStudy (anonymous):

A=5000 * e^(.08*t)

OpenStudy (anonymous):

sorry that is what I meant....how do you solve for t now?

OpenStudy (anonymous):

What's A?

OpenStudy (anonymous):

$7000

OpenStudy (anonymous):

Right, we are trying to grow from 5000 to 7000?

OpenStudy (anonymous):

Yep, make sense, so plug it in in order to to figure out t :)

OpenStudy (anonymous):

7000=5000e^(0.08*t) LN both sides to get LN7000=e^(.08*t) ? Or do you subtract 5000 from 7000 then LN both sides to get LN 2000 = e^(.08*t) ?

OpenStudy (anonymous):

7000=5000e^(0.08*t) -> e^(0.08*t) = 7/5 LN both sides :

OpenStudy (anonymous):

so you divide 7000/5000 and 5000/5000 to separate the e^(.08*t)?

OpenStudy (anonymous):

Always!

OpenStudy (anonymous):

so it should now read LN (7000/5000)=.08*t ?

OpenStudy (anonymous):

ok, got it!!! Thanks for all your help!!!

OpenStudy (anonymous):

The steps always the same :)